INAF LBC+MODS queue observing, 2011 November 26-27

Observer: K. Boutsia, M. Fumana
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: O. Kuhn
Instrument: LBC+MODS

Summary

Clear sky at the beginning of the night, probably photometric .
The seeing at the beginning of the night is high >1.6 and stays between 1.5-1.8'' for most of the night. At the second half we had ~1hr of <1.1'' during which we observe the MODS targtes.
We start with Giallongo because seeing measured on the images is >1.3" and Clementini would be out of specs.
15min lost for the wind measurement - at the moment the procedure started the wind was 15m/s with a max 16.7m/s.
At 7:00UT (midnight MST) we lost preset because of azimuth tracker failure. The TO is worried for possible boggie problem so before going on it was necessary to control the building. 1hr45min lost to this problem.
After the boggie error we continued for a while with LBC because seeing was mediocre (~1.5). When seeing started to get better we switched to MODS.
Lost 2 Red LBC images (2x120sec) because of tracker errors. ~5min lost.

Data Summary

Project Objects Filter ExpTime Notes
Giallongo CL0024+16 g-sloan - r-sloan 3hr the seeing was variable between 1.2 -2.2''- 1hr30min to be repeated
Clementini AndXXV Bbess - Vbess 20min the seeing was ~1.7' - out of specs
Clementini AndXXVII Bbess - Vbess 13min the seeing was ~1.7" - out of spesc - only 2 out of 3 exps because of azimuth tracker problem
STD SA113 Bbess / g-sloan - Vbess / r-sloan   4 positions
Giallongo RXJ0926 g-sloan - r-sloan 1hr the seeing was variable - 30min to be repeated
Donnarruma MAC0717b1 MODS red grating ~1hr 15min the seeing on the guider was <1.1'' - STD + flat +arcs + bias obtained (for calib files look a UT 20111126)

Details

23:24 Turn ON LBC

23:39 Take some test Bias frames to see if everything is working
  • the first 5frames are with the lights ON in the dome.
  • 25Bias_Bino - with the lights OFF

00:39 Taking SkyFlats
  • SkyFlat_BV_pa0 - scale 0.5 : B 47000 / Red 30000
  • SkyFlat_BV_pa180 - sclae 0.8 : B
  • SkyFlat _gr_pa180 - scale 3 - g 30000 / r 20000
  • SkyFlat _gr_pa0 - scale 5 - B 20000 / Red 15000 - too many stars in the flat pointing to zenith - tomorrow we should try a blank field!!

00:58 Slew to Pointing Star ACT0202 for collimation and pointing correction
  • dofpia did not find any pupils in the first interation and exited
  • manually change collimation model.
  • repeat dofpia,/x2,/First
  9889      0      0      0      0      0      0   3.02   3.90     red
  4198   -127     45    -19   -884   -119    -15   2.04   1.71    blue

  -619   -206    302    -41    298     68     26   1.44   0.31     red
  1237   -116   -109   -246    175   -249    -48   1.82   0.54    blue

    3    -79     57   -316    730     55      0   2.35   0.37     red
  -88    -75   -106   -147     87    -47    -14   2.06   0.13    blue
  • DIMM seeing >1.6
  • run lbcrangebal

01:14 Slew to standard star SA113_Bg_Vr

01:17 Filter error - the OB was aborted
  • repeat SA113
  • 4 positions in filters Bbess-g-SLOAB and Vbess-r-SLOAN
  • airmass 1.2
  • red images seem to be out of focus - seeing measured on the red images is double than in the blue side

01:25 Slew to Giallongo_CL0024+16_GR1
  • dofpia
  • we start with Giallongo because seeing measured on the images is >1.2" and Clementini would be out of specs
  • for this program the seeing is within specs
 -2106    360    -30   -517   -882   -485   -100   0.80   1.01     red
   395     49   -247    121   -185     -5     -5   1.59   0.21    blue

  -829    218   -166    641   -721    -57    -76   0.90   0.56     red
 -651     87      7   -302    -59    -19    -46   1.64   0.30    blue

   399    -61   -145   -778    214    -80    -19   0.97   0.41     red
   198   -181     33    270     92    -89    -39   1.38   0.19    blue
  • no hole in the red pupils - stop and do a dofpia,/First
  • manually put Z11+1000 in the red side before proceeding
  • dofpia,/First
  -969    100    -38    651    740   -132    -10   1.22   0.60     red
  -320    124    132    -92    124     11     38   1.45   0.17    blue

  -288    -64     42    105   -434   -104    -37   1.03   0.25     red
 -343     75   -131    204   -108    -53     10   1.49   0.20    blue
  • 3 degrees difference betweein ambient and mirror (-1.6 ambient / 1.5 mirror)

01:38 Giallongo_CL0024+16_GR1
  • seeing is poor - dofpia images look fusy.
  • dimm seeing 1.5''
  • seeing measured on the images: Blue 1.46''/ Red 1.24''
  • seeing measured on the images: Blue 1.2''/ Red 1.1''
  • seeing measured on the images: Blue 1.3''/ Red 1.2''
  • dimm seeing 1.7
  • seeing measured on the images at the end of the ob: ~1.6'' and red starts to go out of focus...

02:18 dofpia
 -3291     28    137     50     94   -356    -52   0.93   1.32     red
-1225   -153   -259    290    892   -752   -100   1.20   0.81    blue

 -270    -10    -90     67    -47   -310    -42   1.05   0.24     red
-856   -126   -165   -105     -6   -153    -35   1.24   0.37    blue

 -346     88    -78    -57     47   -128     -5   0.94   0.18     red
  252   -119   -154     56   -118   -117    -29   1.29   0.17    blue

02:23 Giallongo_CL0024+16_GR2
  • seeing measured on the images: Blue 1.25''/ Red 1.4''
  • seeing measured on the images: ~1.6''
  • seeing measured on the images: Blue 1.7''/ Red 1.55''
  • dimm seeing 2.0''

03:03 dofpia
-1457    -91     12    713    122    -96     13   0.87   0.67     red
-3488    155    316     56    -54    142     85   1.32   1.38    blue

 197   -148    -49   -400   -189    -31     29   1.12   0.23     red
 -163     78   -147   -207     12     87     70   1.75   0.16    blue

03:06 Giallongo_CL0024+16_GR3
  • seeing measured on the images: Blue 1.33''/ Red 1.2''
  • dimm seeing 1.4''
  • seeing measured on the images: Blue 1.35''/ Red 1.25''
  • seeing measured on the images: Blue 1.24''/ Red 1.12''

03:46 dofpia
  
    352   -165    117    -12    315    -10     34   0.89   0.22     red
 -1464   -208    346    274   -240     76     29   1.15   0.62    blue
 
-622    174    144    409    -25      1     38   0.82   0.32     red
 62   -167     41    147    -15    -32     10   1.18   0.12    blue

 124    -94     18   -242     -5    -29     23   0.89   0.15     red
   169     -7    121    -22     14   -131    -11   1.11   0.14    blue

03:51 Giallongo_CL0024+16_GR4
  • dimm seeing >2.0'' but I don't really believe it
  • seeing measured on the images: ~ 1.35''
  • seeing measured on the images: ~ 1.47''
  • seeing measured on the images: Blue 1.6''/ Red 1.45''
  • seeing measured on the images: Blue 1.8''/ Red 1.6''
  • seeing measured on the images: Blue 1.7''/ Red 1.4''

04:31 dofpia
373    252    -72    -25   -206     85     73   1.21   0.21     red
 -2778   -150    -54      5    -11    294    100   1.74   1.11    blue

 -192     72   -125    190     63    174     79   1.76   0.18     red
  157   -106   -144   -201    156    210     76   2.47   0.21    blue

04:35 Giallongo_CL0024+16_GR5
  • seeing measured on the images: Blue 1.8''/ Red 1.6''
  • seeing measured on the images: Blue 1.7''/ Red 1.5''
  • seeing measured on the images: Blue 1.5''/ Red 1.25''
  • most of this OB has seeing between 1.5'' and 1.7''

05:15 Before starting the next OB the TO needs to check the wind
  • at this position the max value is 16.7m/sec

05:26 Slew to Giallongo_CL0024+16_GR6
  • dofpia
  • 15min lost for the wind measurement
 2174    -77     38   -226    -51    136     65   1.80   0.87     red
 -838   -105     14    317   -183    254     65   1.80   0.41    blue

 -999     77    -10    178     22    179     73   1.64   0.43     red
  -249      4     84   -154     52    104     40   1.83   0.16    blue

   -11    -13   -107      6   -142    102     53   1.41   0.12     red
  -153   -103     28     80     -4     11      4   1.40   0.11    blue
  • dimm seeing 2.0''

05:35 Giallongo_CL0024+16_GR6
  • seeing measured on the images: Blue 1.5''/ Red 1.35''
  • seeing measured on the images: ~1.6''
  • image lbcr*.054344 - the sources in the image are elongated - it seems that it was not guiding correctly.
  • seeing measured on the images: Blue 1.9''/ Red 1.7''

06:16 Slew to M31_AndXXV_400s_1 - dofpia
  • dimm seeing >1.6''
 1011    -74     31    -53    -18    -78      1   1.17   0.41     red
 968    101    -96     39   -118   -189    -39   1.27   0.41    blue

  133     54   -112     -5    -21     34     31   1.45   0.10     red
 -82     93    -47   -100     65     70     26   1.61   0.11    blue

06:22 M31_AndXXV_400s_1
  • seeing measured on the images: Blue 1.6''/ Red 1.5''
  • seeing measured on the images: Blue 1.45''/ Red 1.35''
  • seeing measured on the images: ~1.5''

06:45 Slew to M31_AndXXVII_400s_1 - dofpia
  62    -98    -51   -289    106    141     60   1.82   0.19     red
 -666    -27     91    -37    142    290     74   1.92   0.34    blue

-527      0     79    323    -37    -60      4   1.21   0.27     red
  102     92     33     39    -24   -170    -35   1.35   0.14    blue

06:50 M31_AndXXVII_400s_1
  • seeing measured on the images: Blue 1.5''/ Red 1.4''

07:02 We lost tracking - preset cancelled
  • there is a technical problem - it stopped moving in azimuth
  • we stopped the OB and the second image was not saved

07:05 M31_AndXXVII_400s_1 - dofpia
  • the preset was cancelled again
  • the TO is going up the check the boggies

07:28 M31_AndXXVII_400s_1 - just for test, to see if we are able to track
  • in the meantime the TO is checking the building
  • the OB stopped after the first exposure - airmass is out of spec (1.5)

07:40 The TO is going up again to check the boggies
  • after that we switch to MODS

08:42 We can go on - since the seeing is not very good we go on with the Giallongo program

08:44 Slew to Giallongo_RXJ0926_GR1 - dofpia
 
   841    128     13   -848   -508   -187    -43   1.24   0.57     red
 1967    237    175   -668     42    105     -6   1.53   0.84    blue

 -263    -22    110    -86     50     74     12   1.43   0.15     red
  -742    194    -96   -308   -155     76      8   1.53   0.35    blue

  -444      2     75    113     69    110    -15   1.18   0.21     red
  159     67   -125    101     87    -29    -25   1.39   0.13    blue

08:53 Giallongo_RXJ0926_GR1
  • we start at airmass 1.6 ...
  • seeing measured on the images: Blue 1.9''/ Red 1.6''
  • dimm seeing 1.7''
  • seeing measured on the images: Blue 1.5''/ Red 1.4''
  • seeing measured on the images: Blue 1.6''/ Red 1.35''
  • it did not save the 2 last red images ... error with the trackers! - 2 images lost

09:37 dofpia
  1059    -14    -36    -22     19   -145    -18   1.10   0.43     red
   296   -154   -119    172   -117      0      0   2.55   0.18    blue

 -250    146    -94     82    -47   -109    -22   1.17   0.16     red
   271   -221     27   -110    -75    -60     -6   1.38   0.16    blue

09:40 Giallongo_RXJ0926_GR2
  • dimm seeing 1.4''
  • seeing measured on the images: ~1.3''
  • seeing measured on the images: ~1.5'' blue /1.3'' red
  • seeing measured on the images: ~1.3'' blue /1.2'' red
  • dimm seeing 1.4''
  • seeing measured on the images: ~1.1'' blue /1.1'' red
  • seeing measured on the images: ~1.1'' blue /0.9'' red
  • seeing measured on the images: ~'' blue /0.9'' red

10:24 SWITCH TO MODS

10:25 execMODS ~MODSeng/modsScripts/Support/modsWake.pro 0001-0003

10:54 acqMODS mods.7190.201130b1.acq
  • mods1r.20111127.0001.fits - slit image
  • mods1r.20111127.0002.fits - field image
modsAlign mods.7190.201130b1.mms mods1r.20111127.0001.fits mods1r.20111127.0002.fits

Computed Mask Alignment Offset:
   dX:   1.594 arcsec
   dY:  -0.126 arcsec
  dPA: -0.2126 degrees

MODS Offset Command:
   offsetpointing -0.2126  1.594 -0.126 detxy rel
  • mods1r.20111127.0003.fits - throughslit image

11:09 execMODS mods.7190.201130b1.obs 0004-0005
  • guider seeing 0.9''
  • mods1r.20111127.0004.fits - MOS science image
  • seeing on the acquisition image ~0.7''
  • guider <1.1'' - dimm seeing 1.0''
  • mods1r.20111127.0005.fits - MOS science image
  • during the last 600sec of the second image guider seeing ~1.3
  • we are not guiding during the last 10min - 12:14 the last guider image

12:29 Abort after the end of the 2nd exposure
  • we might need to repeat this

12:31 acqMODS standard_star.acq
** ERROR: PRESET IIF error - PresetTelescope result status: Error, no guide star found in guide box after resumeGuiding
  • repeat preset
  • we had some problems with the procedure aliging the slits

12:50 execMODS standard_star.obs 0009
  • the spectrum of the standard looks fine

13:02 acqMODS standard_star1arcsec.acq 0011-0016
  • taking the standard also with the 1arcsec slit - for practice because we are at twilight
  • we must use different scripts for 1arcsec slit and 5arcsec slit ...
  • we centered the star in the slit giving offsets manually

13:13 execMODS standard_star1arcsec.obs
>> partner CALIBRATION
** ERROR: Command 'partner CALIBRATION' timed out after 60 seconds
** Abort, Retry or Ignore? > retry

** standard_star1arcsec.obs aborted with errors at line 3 command 'partner CALIBRATION'
  • repeat exposure

13:15 execMODS standard_star1arcsec.obs 0017
  • we have been guiding until 01:20

13:21 End of night
  • execMODS ~MODSeng/modsScripts/Support/modsSleep.pro

* Seeing reported by GCS during MODS observatioins:
20111127fwhm.jpg

-- KonstantinaBoutsia - 26 Nov 2011
Topic revision: r11 - 23 Feb 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback